HI, i need help with this problem:
A new shipment of 9 VCRs, which will be offered at a very low price, are delivered to a store. Unknow to the manager, 4 of the VCRs are defective.
Determine the probability that at least one of the first 3 VCRs sold is defective. Express your answer to the nearest tenth of one percent.

Answer by stanbon(75887)

A new shipment of 9 VCRs, which will be offered at a very low price, are delivered to a store. Unknow to the manager, 4 of the VCRs are defective.
Determine the probability that at least one of the first 3 VCRs sold is defective. Express your answer to the nearest tenth of one percent.
---------------
P(at least one defective) = 1 - P(none defective)
# of ways to choose 3 that are not defective is 5C3
# of ways to choose 3 of any kind out of nine is 9C3
=1 - 5C3/9C3
=1 - (5/9)(4/8)(3/7)
=1- 0.119047...
=0.8809...

Cheers,
Stan H.